OpenStudy (mollykenney):

factor and apply the zero product property 6x^2−17x=−12

OpenStudy (anonymous):

6x^2 -17x = -12 subtract -12 from both sides 6x^2 - 17x - (-12) = -12 (-12) 6x^2 - 17x + 12 = 0 factor left side of equation (3x - 4)(2x - 3) = 0 set factors equal to 0 3x - 4 = 0 or 2x - 3 = 0 so you have two possiblities for the answer here... x = 4/3 or x = 3/2 what do you think is best?